Jul 23, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. Eastern

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
  4. BCAL HDBCAL log book
  5. Reconstruction and Simulation
  6. Calibration
  7. Any other business


Attendees: Manuel, Elton, Mark, Will, Laura, Shayne (JLab); Andrei (UofR); Sean (NW); Matt (IU)

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
    1. Problem with corrupted base information
      • Problem has been fixed, due to overwriting memory in software
      • Previous monitoring output is suspect and issues re-evaluated with corrected software
      • Test stand is setup, which can be used to extend software development
    2. EPICs migration
      • Hovanes is working on the voltage readout with EPICs. He how has all the necessary equipment. Manuel will work on the GUIs
      • Need this work completed before the fall run
    3. Priorities
      • Check discrete changes in pulse height in a dozen channels. Evaluate if it is due to hardware failures.
      • Set pmt gains assuming the LED light output varies smoothly across the detector.
    4. FCAL Commissioning
      • Draft List for Commissioning
      • The list was discussed and feedback given to plan.
      • Sparcification: Mark mentioned that all readout modes of the FADC have been tested
      • FADC scalers: Should be useful during commissioning. Mark: This mode may require a special event and will look into details.
      • Pi0: Claire used inclusive pi0 production for the simulation of the calibration. Matt is investigating using omega->pi+pi-pi0 as a pi0 source. Could also be used for check reconstruction efficiency.
      • To test the trigger, the idea is to use the LED system.
      • (Sean) How well can the gains be set before beam data? Matt: Hope for +/-20% using the LEDs, which should be sufficient as input to the pi0 gain matching. Timing offsets should be able to be determined quite accurately.
  4. BCAL HDBCAL log book
    1. FADCs
      • Two FADCs were found to have problems in ch 0-7. The problem was traced to very high pedestals (~2000 ch), which were unaffected by the DAC offsets. The first, identified a couple of weeks ago, was replaced. This fixed the problem and the FADC board was checked in the lab and they found some dirt solder joints under one of the chips. The second was identified a few days ago. Will found that the problem was fixed after a soft reboot of the VME controller.
    2. Online monitoring (Mark)
      • Mark has been investigating all the objects starting from low level raw data up through the analysis.
      • Online monitoring histograms have been defined to check them.
      • One problem was found in the simulation, which takes delivers time-over-threshold times as data. This time-walk affected data is ok for the TDC, but not for the FADC, where the algorithm eliminates the time-walk.
      • Need to bring the simulation into alignment with data produced in the hardware. Teagan would be the natural person for this task.
      • Second problem was due to conversion factors in going from raw data to reconstructed data. At the moment the high-level quantities do not have the correct units (and therefore do not make sense). The correct conversions and constants need to be applied to correct this problem. Will and Sean are woking with Mark to correct this problem.
    3. Using muon decays
      • Will has been investigating how to use muon decay electrons as a calibration source for the BCAL
      • Two developments are needed: 1) read out sparcified data so that a long window can be read, and 2) the modules need to be used in self-trigger mode to collect the data.
      • Will has been taking data in pulse integral mode, but still needs the pedestal to be pulled out using JANA. But sparcified data has been taken and multiple pulses are identified in the software. He will need to talk to Dave Lawrence about accessing the pedestal information.
      • (Andrei) Suggested using the top counter alone, or in combination with one of the counters currently below the magnet to enrich cosmic-ray triggers with stopping tracks.
    4. Cosmic-ray rate distribution in BCAL
      • Shayne has plotted the distribution of cosmic-ray hits in the BCAL as a function of module number and it varies quite a bit.
      • To see if the distribution can be understood qualitatively, a simple Monte Carlo is being written to see how the geometry of the counter can affect the distribution.
  5. Reconstruction and Simulation
  6. Calibration
  7. Any other business
    • Note: Meeting abruptly terminated when JLab got disconnected.